# Transfer Learning for Neutrino Scattering: Domain Adaptation with GANs

### v1

@@ Abstract
We utilize transfer learning to extrapolate the physics knowledge encoded in a Generative Adversarial Network (GAN) model trained on synthetic charged-current (CC) neutrino–carbon inclusive scattering data. This base model is adapted to generate CC inclusive scattering events (lepton kinematics only) for neutrino–argon and antineutrino–carbon interactions. Furthermore, we assess the effectiveness of transfer learning in re-optimizing a custom model when new data comes from a different neutrino–nucleus interaction model. Our results demonstrate that transfer learning significantly outperforms training generative models from scratch. To study this, we consider two training data sets: one with 10,000 and another with 100,000 events. The models obtained via transfer learning perform well even with smaller training data. The proposed method provides a promising approach for constructing neutrino scattering event generators in scenarios where experimental data is sparse.
@@ I Introduction
MC generators are often tuned to reproduce experimental observations, relying on adjustable parameters that are fitted using available data [11]. However, this tuning process cannot fully compensate for the fundamental limitations of the underlying models, especially those relying on complex approximations, such as nuclear modeling. Consequently, there is a growing interest in alternative approaches to traditional MC event generation—methods that can learn directly from experimental data and dynamically refine their predictions. Among the most promising directions is the development of artificial intelligence (AI)-driven tools for simulating particle collisions [12, 13, 14].
@@
In this paper, we present results from our development of generative deep learning models designed to simulate neutrino and antineutrino interactions with nuclei. Specifically, we focus on generative adversarial networks (GANs) [15], which have recently gained much interest in high-energy and nuclear physics [16, 17, 18, 19, 20, 21, 22].
@@
Our previous paper [23] presented one of the first applications of GAN models for simulating neutrino collisions with nuclei. Additionally, Refs. [24, 25] provide further examples of using generative deep learning techniques in the simulation of neutrino-nucleus interactions, although in these studies, the method of normalizing flows is employed.
In our previous work, we investigated the properties of GAN models trained to reproduce synthetic inclusive and quasi-elastic charged-current (CC) muon neutrino scattering on carbon nucleus data obtained using the NuWro MC generator. However, our model only generated the kinematics of the charged muon for a given input neutrino energy. In this paper, we explore a technique known as domain adaptation, a form of transfer learning (TL) in which a model trained on one domain (the source domain) is adjusted to perform well on a different but related domain (the target domain). This technique is particularly promising when data in the target domain is scarce.
We previously explored a similar concept in Ref. [26], where a model trained on inclusive electron-carbon scattering data was adapted to predict electron scattering on helium-3, lithium, oxygen, aluminum, calcium, and iron. In this paper, we begin with a baseline model trained on inclusive CC muon neutrino scattering off the carbon nucleus. We then adapt this model to generate predictions for three different processes: muon neutrino scattering off the argon nucleus (relevant for the DUNE experiment), muon-antineutrino scattering off the carbon nucleus, and muon neutrino scattering off the carbon nucleus using data obtained from a different version and configuration of the NuWro generator than that used for training the baseline model.
To evaluate the effectiveness of the TL technique, we compare the results from our domain adaptation models with predictions from models that were trained from scratch. Our findings demonstrate that models employing domain adaptation yield more accurate predictions than those trained from scratch, highlighting the potential of this technique in situations with limited information.
This paper is structured as follows. In Sec. II, key features of the network and its training are explained. In Sec. IV we present the results of the training on multiple neutrino interaction models. Finally, in Sec. V we summarize the paper. In Appendix A, we discuss the properties of the loss function we used in our analysis.
II GAN framework
II.1 Kinematics
@@
The choice of the proxy prime variables is discussed in the Ref. [23]. The values $E_{\nu,min}$ and $E_{\nu,max}$ correspond to minimal and maximal neutrino energy generated by the model, respectively. In this paper, we consider a slightly different value of $E^{-}_{\nu}$ than in the previous study [23]. While the previous boundary prevented the production of non-physical events, they turned out to be too strong for applications discussed in the present studies, namely, $\nu_{\mu}$-argon scattering, or changes in physical parameters in the simulation. Since in experimental data we expect not to follow the same physics as the NuWro simulation, the constraint has been relaxed to $E^{-}_{\nu}=E_{\nu}-4$ MeV.
In our previous studies, we optimized the model using $\nu_{\mu}$-carbon ($^{12}C$) scattering data produced with NuWro ver. 21.09 generator. In the present analysis, we generate data using the same version of NuWro. However, for one discussed case, we consider the newest version of NuWro, ver. 25.03. As we changed the boundary conditions, we repeated the previous analysis and obtained a new model for neutrino-carbon scattering. This model serves as the baseline for subsequent studies.
Figure 1: Generator architecture. The blue block encloses the portion of the network that remains frozen when training with TL.
Figure 2: Discriminator architecture. The blue block encloses the portion of the network that remains frozen when training with TL.
II.2 Networks
We consider the same network architecture as in our previous paper [23] (model for inclusive CC neutrino-carbon scattering). The generator takes as input a random vector of length $50$ (drawn from a Gaussian multivariate distribution) and neutrino energy $E^{\prime}_{\nu}$. Both inputs are processed independently by two distinct dense layers. Then, the output is concatenated and processed by three blocks, as shown in Fig. 1.
To obtain the GAN model, one needs to train the discriminator model, which takes as input the proxy variables $\theta^{\prime}$ and $E_{\mu}$, processed by two independent blocks. The output is then concatenated and transformed. Eventually, the output is transformed by four blocks, as shown in Fig. 2.
The fundamental building block of our network consists of a fully connected (dense) layer, followed by layer normalization [27], an activation function—ReLU[28] for the generator and PReLU[29] for the discriminator, and a dropout layer. We refer to this configuration as Block1. Introducing a skip connection from the input to the output of Block1 yields Block2. When the $E^{\prime}_{\nu}$ information is concatenated to the input of Block1 and subsequently Block2, we obtain Block3 and Block4, respectively. Furthermore, Gaussian noise is injected into both the generator and the discriminator after each layer containing learnable parameters. A comprehensive description of the GAN architecture is provided in Ref.[23].
@@ II.2 Networks
II.3 Loss and metrics
The loss function has two components: one for the discriminator and one for the generator. As in our previous study, the discriminator loss is binary cross-entropy. In the case of the generator loss, we tested three cases, the same choice as in our previous work, the “classical” one initially proposed in Ref. [15], the heuristic (non-saturating) loss also proposed in Ref. [15], and the sum of both. These preliminary studies showed that in the case of our analyses, the last proposal works the most efficiently. As it is not a common choice of the loss, we discuss it in the Appendix A.
@@ II.3 Loss and metrics
Figure 3: Testdata samples $(E^{\prime}_{\mu},\theta^{\prime})$ distributions for $E_{\nu}=500$ MeV. On the top left, we have the original $\nu_{\mu}$-carbon, on the top right is the $\nu_{\mu}$-argon, on the bottom left, $\bar{\nu}_{\mu}$-carbon, and on the bottom right, the alternative $\nu_{\mu}$-carbon samples’ distributions.
Figure 4: Training data $450<E_{\nu}<550$ MeV sub-samples $(E^{\prime}_{\mu},\theta^{\prime})$ distributions. On the left side, we have the sub-sample from a training sample of 100100k events, while on the right side is the sub-sample from a sample of 1010k events.
To evaluate the quality of the model, we consider two metrics: the Mean Averaged Pull (MAP) [32, 18, 23] and the Earth Mover’s Distance (EMD) [33]. In our earlier studies, we used EMD in its two-dimensional form; however, in this analysis, we adapt the MAP to three dimensions (MAP-3D). This adaptation involves creating a histogram of the variables ($E^{\prime}_{\nu}$, $E^{\prime}_{\mu}$, $\theta^{\prime}$) using a uniformly distributed sample of $E_{\nu}$ with a binning of $14\times 14\times 14$.
Typically, the best model is the one that minimizes both metrics, but, in the present analysis, to choose the most optimal model, we evaluate an average of both MAP-3D metrics $\langle\text{MAP}\rangle=\frac{1}{2}(\text{MAP-3D}+\text{MAP-3D w/o tails})$. The selected model minimizes this metric.
III Transfer learning
The main idea of our approach is to use a pre-trained model (baseline model) that was trained on inclusive CC muon neutrino scattering off the carbon target. The new training data are uniformly distributed in the energy range ($300$ MeV, $10$ GeV), similar to our previous analysis.
Using the pre-trained model, we apply the concept of transfer learning [34, 35, 36], assuming that the generator and discriminator have already learned the fundamental patterns underlying neutrino–nucleus interactions. We fine-tune the pre-trained model on smaller training datasets than those used during its initial optimization. This approach leverages the prior knowledge encoded by the model, allowing it to adapt efficiently to new data. A similar strategy was successfully employed in our previous work on modeling inclusive electron scattering data [26].
One of the key advantages of the transfer learning technique is that it reduces the need for large amounts of new data when developing models for related processes. Since the pre-trained model already encodes fundamental relationships underlying neutrino scattering, only a relatively small amount of additional data is required to fine-tune the model to a new but related scattering scenario.
To adapt the pre-trained model to new data, we freeze the first block of both the generator and discriminator, as illustrated in Figs. 1 and 2. The remaining blocks are fine-tuned during the optimization process.
To verify the efficiency of the TL concept, we consider two training data sizes for each process: 100,000 and 10,000 events. Additionally, in each case, we train the model from scratch.
Table 1: Some of the performance metrics of the pre-trained $\nu_{\mu}$-carbon model.
@@ III Transfer learning
Table 2: Performance metrics for generators of CC inclusive $\nu_{\mu}$–argon scattering events. The left and right panels display results for the models trained from scratch and using transfer learning, respectively. The training dataset comprises 10,000 (top table) and 1000,000 (bottom table) events.
@@
III.1 Domain adaptation
We consider three types of inclusive CC scattering processes:
  1. (A)
    $\nu_{\mu}$-argon scattering;
  2. (B)
    $\bar{\nu}_{\mu}$-carbon scattering;
  3. (C)
    $\nu_{\mu}$-carbon scattering but with a different model of interaction than used in the pre-training.
The pre-trained model generates events that follow the distributions predicted by the NuWro ver.21.09 simulation, which employs the spectral function approach to model the carbon (12C) nucleus in the quasielastic (QE) channel [37], and the local Fermi gas model for other interaction channels. The generated events may include a charged muon and one or more nucleons, as well as occasionally pions, resulting from final-state interaction (FSI) effects. Pion production is modeled solely through the $\Delta(1232)$ resonance [38, 39, 40], while contributions from more inelastic processes are modeled using parton distribution functions (PDFs).
In scenario (A), we consider the argon (40Ar) nucleus, described by a spectral function [41] derived from the analysis of Jefferson Lab coincidence electron–argon scattering data [42, 43]. This setup allows us to test the effectiveness of TL when the nuclear target is changed from carbon to a structurally different nucleus.
In Scenario (B), we use the same model for nuclear and hadronic interactions; however, instead of neutrino–carbon interactions, we consider muon antineutrino–carbon interactions.
Scenario (C) addresses a case where the default model used during pre-training encodes interaction dynamics that differ from those in the fine-tuning data. Specifically, the fine-tuning dataset is generated using the latest version of NuWro, ver.25.03, which includes an improved treatment of resonant pion production extended to invariant masses up to $1.9$ GeV. In addition, we deliberately modify key model parameters: the axial mass in the QE channel is increased from $1.030$ GeV to $1.350$ GeV, and the axial mass in the pion production model is changed from $0.94$ GeV to $1.034$ GeV. Finally, the nuclear model is altered by replacing the spectral function with the local Fermi gas model.
The three scenarios differ in their kinematic domains, displaying notable variations in coverage of the relevant phase space, as shown in Fig. 3. This figure presents the $(E^{\prime}_{\mu},\theta^{\prime})$ distributions for 1,000,000 test events in each case: the original test dataset ($\nu_{\mu}$–carbon), and the three test scenarios: $\nu_{\mu}$–argon, $\bar{\nu}_{\mu}$–carbon, and $\nu_{\mu}$–carbon with a modified interaction model. The four datasets exhibit significant differences in their distributions. The most pronounced divergence is observed between the (anti)neutrino–carbon cases and the two involving $\nu_{\mu}$–argon and the altered $\nu_{\mu}$–carbon interaction model. In the latter, a distinct band associated with the QE peak is visible.
Figure 6: Reconstructed hadronic invariant mass $W$-distributions, for inclusive CC $\nu_{\mu}$-argon scattering, generated by NuWro and GAN model trained from scratch and using TL, for the optimization with 10k events. We present the results for $E_{\nu}=500$ MeV (top left), 11 GeV (top right), 33 GeV (bottom left), and 55 GeV (bottom right).
Figure 7: Same as in Fig. 6 but for $C$ $\nu_{\mu}$-argon scattering and the training dataset with 100,000 events.
@@ III.1 Domain adaptation
pRefer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Same as in Fig. 6 but for $C$ $\bar{\nu}_{\mu}$-carbon scattering and the training dataset with 10,000 events.
Figure 9: Same as in Fig. 6 but for $C$ $\bar{\nu}_{\mu}$-carbon scattering and the training dataset with 100,000 events.
@@
Figure 10: Same as in Fig. 6 but for $C$ $\nu_{\mu}$-carbon (alternative interaction) scattering and the training dataset with 10,000 events.

### v2

@@ Abstract
Transfer learning (TL) is used to extrapolate the physics information encoded in a Generative Adversarial Network (GAN) trained on synthetic neutrino-carbon inclusive scattering data to related processes such as neutrino-argon and antineutrino-carbon interactions. We investigate how much of the underlying lepton-nucleus dynamics is shared across different targets and processes. We also assess the effectiveness of TL when training data is obtained from a different neutrino–nucleus interaction model. Our results show that TL not only reproduces key features of lepton kinematics, including the quasielastic and $\Delta$-resonance peaks, but also significantly outperforms generative models trained from scratch. Using data sets of 10,000 and 100,000 events, we find that TL maintains high accuracy even with limited statistics. Our findings demonstrate that TL provides a well-motivated and efficient framework for modeling (anti)neutrino-nucleus interactions and for constructing next-generation neutrino-scattering event generators, particularly valuable when experimental data are sparse.
@@ I Introduction
MC generators are often tuned to reproduce experimental observations, relying on adjustable parameters fitted to available data [11]. However, this tuning process cannot fully compensate for the fundamental limitations of the underlying models, especially those relying on complex approximations, such as nuclear modeling. Consequently, there is a growing interest in alternative approaches to traditional MC event generation—methods that can learn directly from experimental or synthetic data and dynamically refine their predictions. Among the most promising directions is the development of artificial intelligence (AI)-driven tools for simulating particle collisions [12, 13, 14].
@@
In this paper, we present results from our development of generative deep learning models to simulate neutrino and antineutrino interactions with nuclei. Specifically, we focus on generative adversarial networks (GANs) [15], which have recently attracted much interest in high-energy and nuclear physics [16, 17, 18, 19, 20, 21, 22].
@@
Our previous paper [23] presented one of the first applications of GAN models for simulating neutrino collisions with nuclei. Additionally, Refs. [24, 25] provide further examples of using generative deep-learning techniques in the simulation of neutrino–nucleus interactions, although in those studies, a method of normalizing flows was employed.
The main objective of this study is to examine the generalization ability of deep generative models, following the idea introduced in our earlier work [26]. There, we showed that a model trained on inclusive electron–carbon scattering data could be adapted via transfer learning (TL) to predict cross sections for electron scattering on helium-3, lithium, oxygen, aluminum, calcium, and iron. This suggested that the network had captured universal features of lepton–nucleus dynamics, such as kinematic constraints and the relative position and strengths of quasielastic and $\Delta(1232)$ resonance peaks.
Here, we extend that concept to neutrino interactions. We employ a pre-trained GAN model [23], trained on synthetic inclusive charged-current (CC) muon-neutrino scattering data obtained from the NuWro MC generator, as a baseline model encoding nuclear relationships. The model generates lepton kinematics in the quasielastic, $\Delta(1232)$ resonance and inelastic regions for a given input neutrino energy.
We hypothesize that this model retains generic relationships common to CC neutrino scattering processes, independent of the target nucleus or probe type. To test this, we apply domain adaptation, a form of TL in which a model trained in one domain (the source) is fine-tuned to operate effectively in another, related domain (the target). This allows us to probe how much of the encoded physics can be reused when data are limited or originate from a different generator configuration. The baseline model is adapted to generate predictions for three processes: (A) muon-neutrino scattering off an argon nucleus; (B) muon-antineutrino scattering off a carbon nucleus; and (C) muon-neutrino scattering off a carbon nucleus, using data from an alternative version and configuration of NuWro.
By comparing the domain-adapted and from-scratch models, we demonstrate that TL preserves the essential quasielastic and $\Delta(1232)$-resonance structure of the interactions and achieves high accuracy even with reduced training statistics. The GAN primarily encodes universal kinematic constraints and the overall quasielastic (QE) and $\Delta$-resonance pattern, while target-dependent spectral-function details are corrected by fine-tuning. These findings suggest that TL is not only a computationally efficient [^1] A computational efficiency of TL in classification of neutrino interaction processes was studied in Ref. [27]. method but also a physically interpretable tool for constructing (anti)neutrino–nucleus scattering models, enabling insight into which aspects of lepton–hadron dynamics are universal and which require re-optimization.
This paper is structured as follows. In Sec. II, we discuss the idea of transfer learning and its applications to building (anti)neutrino-scattering models. In Sec. III, key features of the GAN models and training procedures are explained. In Sec. IV, we present the numerical results. Finally, in Sec. V we summarize the paper. The paper is accompanied by supplemental material Sections S1, S2, and S3, which includes figures showing the performance of the obtained models and a detailed description of the loss functions used in the numerical analysis.
II Neutrino interactions from transfer learning
II.1 Transfer learning
The TL or knowledge transfer originally emerged from pedagogical and psychological research. [28]. It is founded on the idea that skills acquired in one task, such as playing the piano, can be applied to solve different tasks, like mathematical problems. Both types of challenges are addressed using the same set of neurons, which can either be used directly for both tasks or be slightly modified to tackle different kinds of problems.
TL has emerged as a powerful technique in deep learning [29, 30, 31, 32]. Models that are initially trained for one type of problem can later be adapted, typically through fine-tuning, to address a different problem. The fundamental idea is that the initial layers of a neural network, once adapted to a specific task, extract universal features from the data that can be applied almost directly to related problems.
Our goal is to showcase the effectiveness of TL in simulating (anti)neutrino-nucleus interactions. Specifically, we aim to begin with a GAN trained on synthetic data generated by the theoretical framework, particularly the NuWro MC generator. The predictions made by this generator align with reality to some extent, like any other theoretically based model. Therefore, the GAN model trained on NuWro generated data will serve as a reference for creating more realistic simulators.
In our previous studies [26], using the TL technique, we showed that the inclusive electron-carbon interaction model can be adapted with a small amount of information to predict electron scattering off targets ranging from helium-3 to iron. In this paper, we demonstrate that, from a GAN simulation of neutrino-carbon scattering, one can obtain models for neutrino-argon, antineutrino-carbon, and neutrino-carbon (with a different nuclear model than the reference) scattering.
Similarly to Ref. [26], we assume that GAN’s generator and discriminator for neutrino-carbon interactions, obtained in our previous work [23], have already learned the fundamental patterns underlying neutrino–nucleus interactions, and we will fine-tune this pre-trained model to obtain models for related processes.
One of the key advantages of the TL technique is that it reduces the need for large amounts of new data when developing models for related processes. Since the pre-trained model already encodes the fundamental relationships underlying neutrino scattering, only a relatively small amount of additional data is needed to fine-tune it. This property makes TL particularly attractive for neutrino–nucleus interactions, where experimental data are often sparse.
Figure 1: Test data samples $(E^{\prime}_{\mu},\theta^{\prime})$ distributions for $E_{\nu}=500$ MeV. On the top left, we show the original MC $\nu_{\mu}$-carbon scattering; on the top right, the $\nu_{\mu}$-argon scattering; on the bottom left, the $\bar{\nu}_{\mu}$-carbon scattering; and on the bottom right, the distributions of the alternative $\nu_{\mu}$-carbon samples.
II.2 Domain adaptation
As mentioned above, we consider the GAN model generating muon kinematical variables for the inclusive CC muon neutrino-carbon scattering. This model will be fine-tuned to obtain a description of three types of inclusive CC scattering processes:
  1. (A)
    $\nu_{\mu}$-argon scattering;
  2. (B)
    $\bar{\nu}_{\mu}$-carbon scattering;
  3. (C)
    $\nu_{\mu}$-carbon scattering with a different model of interaction than one used in the pre-training.
All models assume a similar (uniform) initial spectrum of energy for (anti)neutrinos, and the target space encompasses configurations of kinematic variables that characterize the charged muon. Because the pre-trained model is trained on data from a single process (source domain), the resulting models are expected to generalize to related processes defined in a distinct domain. Consequently, the TL approach employed in this paper can be regarded as an instance of domain adaptation.
The pre-trained model generates events as predicted by the NuWro ver.21.09, which employs the spectral function approach to model the carbon (12C) nucleus in the QE channel [33], and the local Fermi gas (LFG) model for other interaction channels. The generated events may include a charged muon and one or more nucleons, as well as occasionally pions, resulting from final-state interaction (FSI) effects. Pion production is modeled mostly through the $\Delta(1232)$ resonance excitation [34, 35, 36], while contributions from more inelastic processes are modeled using parton distribution functions (PDFs).
In scenario (A), we consider the argon (40Ar) nucleus, described by a spectral function (SF) [37] derived from the analysis of Jefferson Lab coincidence electron–argon scattering data [38, 39]. This setup allows us to test the effectiveness of TL when the nuclear target is changed from carbon to a structurally different nucleus. Note that the NuWro SF model for carbon includes an optical potential, which is not present in the argon model [^2] The latest version of NuWro keeps optical potential for both targets..
In Scenario (B), we use the same model for nuclear and hadronic interactions; however, instead of neutrino–carbon interactions, we consider muon antineutrino–carbon interactions. From a theoretical point of view, on the three-level, the main difference between neutrino and antineutrino interactions is the sign between vector and axial contributions in the hadronic transition current. It results in the redistribution of the differential cross sections and the suppression of the total cross section with respect to the neutrino scattering cross section.
Scenario (C) addresses a case where the default model used during pre-training encodes interaction dynamics that differ from those in the fine-tuning data. Specifically, the fine-tuning dataset is generated using the version of NuWro 25.03, which includes an improved treatment of resonant pion production extended to invariant masses up to $1.9$ GeV. In addition, we deliberately modify key model theoretical parameters: the axial mass in the QE channel is increased from $1.03$ GeV to $1.350$ GeV, and the axial mass in the pion production model is changed from $0.94$ GeV to $1.034$ GeV. Finally, the nuclear model is altered by replacing the spectral function with the LFG model.
The three scenarios differ in their kinematic domains, displaying notable variations in coverage of the relevant phase space, as shown in Fig. 1. This figure presents the $(E^{\prime}_{\mu},\theta^{\prime})$ distributions, see Sec. III.1, for 1,000,000 test events, for $E_{\nu}=500$ MeV, in each case: the original test dataset ($\nu_{\mu}$–carbon), and the three test scenarios: $\nu_{\mu}$–argon, $\bar{\nu}_{\mu}$–carbon, and $\nu_{\mu}$–carbon with a modified interaction model [^3] Note that in the supplement we show also $(E_{\mu},\theta)$ distributions for $E_{\nu}=500$ MeV for neutrino scattering with carbon and argon.. All the distributions are strongly dominated by QE events. The four datasets exhibit significant differences in their distributions. The most pronounced divergence is observed between the (anti)neutrino–carbon cases and the two involving $\nu_{\mu}$–argon and the altered $\nu_{\mu}$–carbon interaction model. In the latter, a distinct band associated with the QE peak is visible. The differences observed between neutrino–carbon and neutrino–argon event distributions arise from the distinct spectral functions of the two nuclei. Argon, being much heavier and more complex than carbon, with more occupied shells, more nucleons, and a stronger Coulomb potential, has a spectral function characterized by larger and more widely distributed removal energies, as well as a different (and potentially stronger) high-momentum component. There is also an important effect originating from the optical potential included in the carbon spectral function. Consequently, the QE peak in argon is broader than that observed for carbon targets (see Fig. S1 in the supplement). In the antineutrino–carbon event distribution, the QE peak is less pronounced than in the neutrino case. Finally, in the alternative carbon model using the LFG, the QE peak becomes noticeably sharper.
III GAN framework
III.1 Kinematics
Figure 2: Generator architecture. The blue block encloses the portion of the network that remains frozen when training with TL. The output of the network is given by the proxy vector Eq. 4. Similar to Ref. [23], the skip connections are applied but not shown in the graph above.
Figure 3: Discriminator architecture. The blue block encloses the portion of the network that remains frozen when training with TL. The network takes for the input $\mathbf{v}$ (the proxy variable vector defined by Eq. 4) and proxy neutrino energy $E^{\prime}_{\nu}$. Similar to Ref. [23], the skip connections are applied but not shown in the graph above.
@@ III.1 Kinematics
Note that our research plan involves generating comprehensive event information, including the hadronic components. However, creating all final-state particles is complex due to the variability in the number of final particles and the need to ensure charge, energy, and momentum conservation. For this reason, we have chosen first to investigate inclusive contributions to test models and various solutions.
@@
The choice of the proxy variables is discussed in the Ref. [23]. The values $E_{\nu,min}$ and $E_{\nu,max}$ correspond to minimal and maximal neutrino energy generated by the model. In this paper, we consider a slightly different value of $E^{-}_{\nu}$ than in the previous study [23]. While the previously introduced boundaries prevented the production of non-physical events, they turned out to be too strong for the applications discussed in the present study. Since in experimental data we expect not to follow the same physics as the NuWro simulation, the constraint has been relaxed to $E^{-}_{\nu}=E_{\nu}-4$ MeV.
In our previous studies, we optimized the model using $\nu_{\mu}$-carbon (12C) scattering data produced with NuWro ver. 21.09 generator. In the present analysis, we generate data using the same version of NuWro. However, for one discussed case, we consider the NuWro ver. 25.03. As we changed the boundary conditions, we repeated the previous analysis and obtained a new model for neutrino-carbon scattering. This model serves as the baseline for subsequent studies.
III.2 Networks
We consider the same network architecture as in our previous paper [23] (model for inclusive CC neutrino-carbon scattering). The generator takes as input a random vector of length $50$ (drawn from a Gaussian multivariate distribution) and neutrino energy $E^{\prime}_{\nu}$. Both inputs are processed independently by two distinct dense layers. Then, the output is concatenated and processed by three blocks, as shown in Fig. 2. The output of the generator is given by a last dense layer that produces the proxy vector
$\mathbf{v}=(\theta^{\prime},{E^{\prime}}_{\mu}).$ (4)
To obtain the GAN model, one needs to train the discriminator model, which takes as input the proxy variables $\mathbf{v}$ and neutrino energy $E^{\prime}_{\nu}$, which are concatenated and processed by two independent blocks. The output is then concatenated and transformed. Eventually, the output is transformed by four blocks, as shown in Fig. 3. Finally, the discriminator uses dense layer with single output value.
The fundamental building block of our network consists of a fully connected (dense) layer, followed by layer normalization [40], an activation function—ReLU[41] for the generator and PReLU[42] for the discriminator, and a dropout layer. We refer to this configuration as Block1. Introducing a skip connection from the input to the output of Block1 yields Block2 (the latter is not used in the models of the present study). When the $E^{\prime}_{\nu}$ information is concatenated to the input of Block1 and subsequently Block2, we obtain Block3 and Block4, respectively. Furthermore, Gaussian noise is injected into both the generator and the discriminator after each layer containing learnable parameters. A comprehensive description of the GAN architecture is provided in Ref.[23].
@@ III.2 Networks
To adapt the pre-trained model to new data, we freeze the first block of both the generator and discriminator, as illustrated in Figs. 2 and 3. The remaining blocks are fine-tuned during the optimization process.
III.3 Loss and metrics
The loss function has two components: one for the discriminator and one for the generator. As in our previous study, the discriminator loss is binary cross-entropy. In the case of generator loss, we tested three cases: the same choices as in our previous work, the “classical” one initially proposed in Ref. [15], the heuristic (non-saturating) loss also proposed in Ref. [15], and the sum of both. These preliminary studies showed that the last proposal is the most efficient. As it is not a common choice of the loss, we discuss its properties in the supplement section S3.
@@ III.3 Loss and metrics
Table 1: Performance metrics of the pre-trained $\nu_{\mu}$-carbon model.
@@
To evaluate the quality of the model, we consider two metrics: the Mean Averaged Pull (MAP) [45, 18, 23] and the Earth Mover’s Distance (EMD) [46]. In our earlier studies, we used MAP in its two-dimensional form; however, in this analysis, we adapt the MAP to three dimensions (MAP-3D). This adaptation involves creating a histogram of the variables ($E^{\prime}_{\nu}$, $E^{\prime}_{\mu}$, $\theta^{\prime}$) using a uniformly distributed sample of $E_{\nu}$ with a binning of $14\times 14\times 14$ [^4] The number of bins in the MAP-3D calculation is approximately similar to the number of $50\times 50$ bins used for the 2-dimensional version in the previous version..
Typically, the best model is the one that minimizes both metrics, but, in the present analysis, to choose the optimal model, we evaluate an average of both MAP-3D metrics $\langle\text{MAP}\rangle=\frac{1}{2}(\text{MAP-3D}+\text{MAP-3D w/o tails})$.
Figure 4: Training data $450<E_{\nu}<550$ MeV sub-samples $(E^{\prime}_{\mu},\theta^{\prime})$ distributions. On the top, we have the sub-sample from a training sample of 100100k events, while on the bottom is the sub-sample from a sample of 1010k events.
Figure 6: Reconstructed hadronic invariant mass $W$-distributions, for $E_{\nu}=1$ GeV and for inclusive CC $\nu_{\mu}$-argon scattering, generated by NuWro and GAN model trained from scratch and using TL, for the optimization with 10,000 (top) and 100,000 (bottom) events (top).
Table 2: Performance metrics for generators of $C$ inclusive $\nu_{\mu}$–argon scattering events. The left and right panels display results for the models trained from scratch and using transfer learning, respectively. The training dataset comprises 10,000 (top table) and 100,000 (bottom table) events.
@@
pRefer to caption
Figure 7: Same as in Fig. 6 but for $C$ $\bar{\nu}_{\mu}$-carbon scattering and $E_{\nu}=500$ MeV.
@@
Figure 8: Same as in Fig. 6 but for $C$ $\nu_{\mu}$-carbon, $E_{\nu}=1$ GeV (alternative interaction) scattering and the training dataset with 10,000 events.